# TFB: Towards Comprehensive and Fair Benchmarking of Time Series Forecasting Methods

### v2

@@ 4.3.2. Evaluation metrics
To achieve an extensive evaluation of forecasting performance, we employ eight error metrics, namely the Mean Absolute Error (MAE), Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), Symmetric Mean Absolute Percentage Error (SMAPE), Weighted Absolute Percent Error (WAPE), Mean Absolute Scaled Error (MASE) (Hyndman and Koehler, 2006), and Modified Symmetric Mean Absolute Percentage Error (MSMAPE) (Suilin, 2017), which are defined in Equations  7– 14.
@@ 5.1. Experimental Setup
5.1.1. Datasets
We utilized all the datasets included in TFB, encompassing 25 multivariate datasets and 8,068 univariate time series. The detailed data splits for the multivariate datasets are given in Table 5, and the forecasting horizons $F$ of the univariate datasets are given in Table 4.
5.1.2. Comparison methods
To reduce the stereotype bias on comparison methods, all methods mentioned in Section 4.2 are included for a comprehensive comparison.
@@ 5.2.3. Performance on different characteristics
These observations indicate that various deep learning methods have varying abilities on datasets with different characteristics.
@@
ZafNoo 96 0.426 0.444 0.419 0.432 0.441 0.475 0.473 0.541 0.420 0.442 0.411 0.434 0.410 0.446 0.421 0.442 0.424 0.479 0.474 0.533 0.411 0.451 1.274 3.928 1.2e+160 inf 1.4e+229 inf
192 0.456 0.498 0.449 0.479 0.479 0.544 0.575 0.708 0.448 0.493 0.444 0.484 0.447 0.503 0.455 0.493 0.446 0.491 0.476 0.546 0.448 0.508 1.271 3.858 nan nan nan nan
336 0.480 0.530 0.469 0.521 0.517 0.595 0.661 0.851 0.467 0.523 0.464 0.518 0.470 0.544 0.455 0.504 0.479 0.551 0.507 0.624 0.471 0.549 1.238 3.586 nan nan nan nan
720 0.499 0.574 0.483 0.543 0.560 0.697 0.699 0.876 0.494 0.563 0.486 0.548 0.504 0.595 0.476 0.540 0.511 0.627 0.508 0.599 0.502 0.586 1.169 3.053 nan nan nan nan
CzeLan 96 0.251 0.183 0.443 0.581 0.310 0.231 0.305 0.250 0.519 0.818 0.289 0.211 0.229 0.178 0.275 0.194 0.237 0.176 0.553 0.519 0.232 0.180 1.312 3.143 1.6e+75 4.2e+152 0.351 0.302
192 0.271 0.208 0.503 0.705 0.339 0.268 0.337 0.295 0.589 0.962 0.323 0.252 0.252 0.210 0.322 0.254 0.279 0.215 0.559 0.539 0.255 0.212 1.233 2.963 2.6e+151 inf 0.401 0.368
336 0.302 0.243 0.596 0.971 0.361 0.298 0.361 0.335 0.659 1.161 0.366 0.317 0.280 0.243 0.362 0.314 0.288 0.224 0.692 0.966 0.281 0.243 1.243 2.985 3.6e+265 inf 0.454 0.443
720 0.335 0.273 0.762 1.566 0.446 0.416 0.416 0.384 0.741 1.496 0.392 0.358 0.317 0.284 0.396 0.354 0.337 0.282 0.848 1.143 0.305 0.268 1.257 3.020 nan nan 0.524 0.544
Covid-19 24 0.042 1.052 2.297 1768.432 0.177 2.125 0.079 2.642 2.395 1768.993 0.240 9.587 0.070 1.139 0.438 33.908 0.064 2.231 55.555 6274.889 0.047 1.183 3.601 1711.553 205.153 4.4e+6 7.056 5.8e+4
36 0.051 1.398 2.394 1770.695 0.177 2.463 0.087 3.064 2.432 1769.541 0.173 5.333 0.091 1.582 0.494 46.864 0.081 2.530 72.837 1.1e+4 0.059 1.470 3.600 1711.515 224.069 5.2e+6 10.333 9.3e+4
48 0.060 1.814 2.465 1773.221 0.192 2.847 0.097 3.698 2.481 1770.224 0.246 9.598 0.099 1.932 0.566 46.563 0.079 2.561 84.309 1.3e+4 0.069 1.859 3.600 1711.519 407.660 3.3e+7 11.473 8.5e+4
60 0.068 2.265 2.427 1772.835 0.212 3.275 0.106 4.232 2.533 1773.691 0.297 7.778 0.127 2.682 0.711 94.629 0.098 3.675 93.938 1.6e+4 0.078 2.278 3.600 1711.542 616.234 6.1e+7 15.914 9.6e+4
NASDAQ 24 0.567 0.649 0.745 1.149 0.547 0.627 0.744 1.110 1.330 2.727 0.666 0.830 0.522 0.566 0.630 0.776 0.539 0.563 1.554 3.153 0.645 0.767 0.608 0.752 0.705 1.017 0.462 0.516
36 0.682 0.821 0.885 1.414 0.659 0.885 0.895 1.399 1.534 3.387 0.862 1.356 0.656 0.827 0.832 1.325 0.687 0.905 1.693 3.853 0.835 1.379 0.726 1.062 0.944 1.701 0.653 0.920
48 0.793 1.169 1.136 2.108 0.786 1.139 0.864 1.296 1.551 3.412 0.990 1.817 0.757 1.106 0.977 1.788 0.783 1.218 1.637 3.582 0.829 1.179 0.790 1.263 1.131 2.411 0.805 1.366
60 0.828 1.268 1.230 2.336 0.783 1.251 0.834 1.171 1.528 3.287 1.037 2.011 0.792 1.204 1.099 2.230 0.781 1.298 2.017 4.984 0.853 1.303 0.814 1.328 1.264 3.042 0.879 1.619
@@ Acknowledgements.
This work was partially supported by National Natural Science Foundation of China (62372179), Independent Research Fund Denmark under agreement 8048-00038B, and Huawei Cloud Algorithm Innovation Lab.

### v3

@@ 4.3.2. Evaluation metrics
To achieve an extensive evaluation of forecasting performance, we employ eight error metrics, namely the Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), Mean Squared Error (MSE), Symmetric Mean Absolute Percentage Error (SMAPE), Root Mean Squared Error (RMSE), Weighted Absolute Percent Error (WAPE), Modified Symmetric Mean Absolute Percentage Error (MSMAPE) (Suilin, 2017), and Mean Absolute Scaled Error (MASE) (Hyndman and Koehler, 2006), which are defined in Equations  7– 14.
@@ 5.1. Experimental Setup
5.1.1. Datasets and Comparison methods
We include all the datasets included in TFB and all methods mentioned in Section 4.2.
@@ 5.2.3. Performance on different characteristics
ZafNoo 96 0.426 0.444 0.419 0.432 0.441 0.475 0.473 0.541 0.420 0.442 0.411 0.434 0.410 0.446 0.421 0.442 0.424 0.479 0.474 0.533 0.411 0.451 1.274 3.928 0.401 0.412 1.4e+229 inf
192 0.456 0.498 0.449 0.479 0.479 0.544 0.575 0.708 0.448 0.493 0.444 0.484 0.447 0.503 0.455 0.493 0.446 0.491 0.476 0.546 0.448 0.508 1.271 3.858 0.438 0.472 nan nan
336 0.480 0.530 0.469 0.521 0.517 0.595 0.661 0.851 0.467 0.523 0.464 0.518 0.470 0.544 0.455 0.504 0.479 0.551 0.507 0.624 0.471 0.549 1.238 3.586 0.454 0.506 nan nan
720 0.499 0.574 0.483 0.543 0.560 0.697 0.699 0.876 0.494 0.563 0.486 0.548 0.504 0.595 0.476 0.540 0.511 0.627 0.508 0.599 0.502 0.586 1.169 3.053 0.476 0.554 nan nan
CzeLan 96 0.251 0.183 0.443 0.581 0.310 0.231 0.305 0.250 0.519 0.818 0.289 0.211 0.229 0.178 0.275 0.194 0.237 0.176 0.553 0.519 0.232 0.180 1.312 3.143 0.262 0.184 0.351 0.302
192 0.271 0.208 0.503 0.705 0.339 0.268 0.337 0.295 0.589 0.962 0.323 0.252 0.252 0.210 0.322 0.254 0.279 0.215 0.559 0.539 0.255 0.212 1.233 2.963 0.300 0.222 0.401 0.368
336 0.302 0.243 0.596 0.971 0.361 0.298 0.361 0.335 0.659 1.161 0.366 0.317 0.280 0.243 0.362 0.314 0.288 0.224 0.692 0.966 0.281 0.243 1.243 2.985 0.357 0.287 0.454 0.443
720 0.335 0.273 0.762 1.566 0.446 0.416 0.416 0.384 0.741 1.496 0.392 0.358 0.317 0.284 0.396 0.354 0.337 0.282 0.848 1.143 0.305 0.268 1.257 3.020 0.445 0.405 0.524 0.544
Covid-19 24 0.042 1.052 2.297 1768.432 0.177 2.125 0.079 2.642 2.395 1768.993 0.240 9.587 0.070 1.139 0.438 33.908 0.064 2.231 55.555 6274.889 0.047 1.183 3.601 1711.553 0.727 80.511 7.056 5.8e+4
36 0.051 1.398 2.394 1770.695 0.177 2.463 0.087 3.064 2.432 1769.541 0.173 5.333 0.091 1.582 0.494 46.864 0.081 2.530 72.837 1.1e+4 0.059 1.470 3.600 1711.515 0.644 51.959 10.333 9.3e+4
48 0.060 1.814 2.465 1773.221 0.192 2.847 0.097 3.698 2.481 1770.224 0.246 9.598 0.099 1.932 0.566 46.563 0.079 2.561 84.309 1.3e+4 0.069 1.859 3.600 1711.519 0.635 52.277 11.473 8.5e+4
60 0.068 2.265 2.427 1772.835 0.212 3.275 0.106 4.232 2.533 1773.691 0.297 7.778 0.127 2.682 0.711 94.629 0.098 3.675 93.938 1.6e+4 0.078 2.278 3.600 1711.542 0.737 102.415 15.914 9.6e+4
NASDAQ 24 0.567 0.649 0.745 1.149 0.547 0.627 0.744 1.110 1.330 2.727 0.666 0.830 0.522 0.566 0.630 0.776 0.539 0.563 1.554 3.153 0.645 0.767 0.608 0.752 0.616 0.774 0.462 0.516
36 0.682 0.821 0.885 1.414 0.659 0.885 0.895 1.399 1.534 3.387 0.862 1.356 0.656 0.827 0.832 1.325 0.687 0.905 1.693 3.853 0.835 1.379 0.726 1.062 0.803 1.223 0.653 0.920
48 0.793 1.169 1.136 2.108 0.786 1.139 0.864 1.296 1.551 3.412 0.990 1.817 0.757 1.106 0.977 1.788 0.783 1.218 1.637 3.582 0.829 1.179 0.790 1.263 0.893 1.540 0.805 1.366
60 0.828 1.268 1.230 2.336 0.783 1.251 0.834 1.171 1.528 3.287 1.037 2.011 0.792 1.204 1.099 2.230 0.781 1.298 2.017 4.984 0.853 1.303 0.814 1.328 0.964 1.753 0.879 1.619
@@ Acknowledgements.
This work was partially supported by National Natural Science Foundation of China (62372179), Independent Research Fund Denmark under agreement 8048-00038B, and Huawei Cloud Algorithm Innovation Lab. Jilin Hu is the corresponding author of the work.